OpenStudy (anonymous):

Find sin(pi/3 - pi/4)

OpenStudy (mertsj):

\[\sin (x-y)=\sin x \cos y-\cos x \sin y\]

OpenStudy (mathmale):

For reference, Baby Doll: sin pi/3 = Sqrt(30/2 cos pi/3 = 1/2 Can you provide the values for sin pi/4 and cos pi/4?